OpenStudy (anonymous):

if 294 grams of FeS2 is allowed ti react with 176 grams of O2 according to the following equation how many grams of Fe2O3 are produced? FeS2 + O2 = Fe2O3 + SO2

OpenStudy (koikkara):

@ayat7654 Now, it's your turn... Solve it, then i'll help you when u get stuck.

OpenStudy (anonymous):

@koikkara :) I hope it's 👍🏻

OpenStudy (koikkara):

Nope, wrong... You hav to balance the eqn first and then start with it.... \(FeS2 + O2 = Fe2O3 + SO2 \) Balanced Eqn: \(4 FeS2 + 11 O2 = 2 Fe2O3 + 8 SO2\) Substituting values, \(294g.FeS2× 1mol.FeS2/119.99× 11mols .O2/4mols .FeS2\) \(= 6.738mol .O2\) \(176gO2×1mol O2/32gO2× 4mols FeS2/11mol FeS2\)\(= 2mols FeS2 \) select, molecule with the lowest amount (Limiting Reagent) \(2molsFeS2×2molsFe2O3/4molsFeS2×159.7gF\)... \(159.7g. Fe2O3( grams) produced\) @ayat7654

OpenStudy (anonymous):

@koikkara

OpenStudy (anonymous):

@koikkara Oh..by mistake But thanks